OpenStudy (anonymous):

Help asap been sick really need to catch up Which of the following numbers is prime? 133 143 153 163 Question 2. Determine the prime factorization of 88. (Points : 1) 2 • 32• 5 22• 3 • 5 23• 11 22• 11 Question 3. 3. Determine the prime factorization of 616. (Points : 1) 23 • 7 • 11 22 • 7 • 13 23 • 3 • 17 22 • 53 Question 4. 4. Find the greatest common factor of the pair of numbers. 198 and 330 99 6

OpenStudy (anonymous):

. Find the greatest common factor of the pair of numbers. 198 and 330 (Points : 1) 99 66 30 11 Question 5. 5. Find the greatest common factor of the pair of numbers. 333 and 441 (Points : 1) 3 6 9 27

ganeshie8 (ganeshie8):

for question #1 : 133 = 7 * 19, so 133 is NOT a prime number 143 = 11 * 13, so 143 is NOT a prime number 153 = 9 * 17, so 153 is NOT a prime number 163 = 1 * 163, so 163 is a prime number \(\large \color{red}{\checkmark}\)

ganeshie8 (ganeshie8):

does that make some sense ? basically if a number has no divisors other than 1 and itself, then its a prime number

OpenStudy (anonymous):

thanks:)